OpenStudy (anonymous):

An angle measures 2 degrees more than 3 times its complement. Find the measure of its complement

OpenStudy (lgbasallote):

complement is 90 - x. So your equation is x =2 + 3(90-x) which becomes 2 + 270 - 3x = x. So, 4x = 272. x = 68. So its complement is equal to 22.